OpenStudy (anonymous):

how long does it take for a deposit of $P to double to 10% interest compounded continuously?

Directrix (directrix):

A = P e ^(rt) is the formula for continuous compounding of interest Let P = amount invested and 2P the desired amount after t years at 10% interest (2P) = P e ^ (.1t) 2 = 1 * e ^ (.1t) 2 = e ^ (.1t) ln(2) = ln (e ^ (.1t)) ln (2) = .1t * ln(e) ln (2) = .1t * 1 ln (2) = .1t t = [ ln(2) ] / (.1) t = 6.9314 years for principal investment to double

OpenStudy (paxpolaris):

\[\Large A=P \cdot e^{rt}\] r=10/100 =0.1 solve for 't' when A=2P

OpenStudy (anonymous):

simple just divide 72 by number of years

OpenStudy (anonymous):

it is called as rule of 72

OpenStudy (anonymous):

alot of time